# Safety in Large Reasoning Models: A Survey

### v1

@@ Abstract
Large Reasoning Models (LRMs) have exhibited extraordinary prowess in tasks like mathematics and coding, leveraging their advanced reasoning capabilities. Nevertheless, as these capabilities progress, significant concerns regarding their vulnerabilities and safety have arisen, which can pose challenges to their deployment and application in real-world settings. This paper presents a comprehensive survey of LRMs, meticulously exploring and summarizing the newly emerged safety risks, attacks, and defense strategies. By organizing these elements into a detailed taxonomy, this work aims to offer a clear and structured understanding of the current safety landscape of LRMs, facilitating future research and development to enhance the security and reliability of these powerful models.
@@ 1 Introduction
Large Language Models (LLMs) (Meta, 2024; Qwen et al., 2025) have achieved remarkable proficiency across tasks ranging from open-domain conversation to program synthesis. Central to their utility is reasoning: the ability to derive logically coherent conclusions by chaining together intermediate inferences.
@@
The recent release of OpenAI’s o1 series (OpenAI, 2024) marks the emergence of Large Reasoning Models (LRMs), which are explicitly trained to produce richly formatted, human-readable reasoning traces. Notable examples include DeepSeek-R1 (DeepSeek-AI et al., 2025), Kimi-1.5 (Team et al., 2025), and QwQ (Team, 2024b), all of which leverage reinforcement learning to refine their deduction processes. LRMs now set new benchmarks in mathematical problem solving (Lightman et al., 2023), closed-book question answering (Rein et al., 2024), and code generation (Jain et al., 2024).
As LRMs become increasingly integrated into high-stakes domains—from scientific research to autonomous decision support—it is vital to rigorously assess their safety, robustness, and alignment. Despite there are already a lot of surveys providing on LLM safety (Huang et al., 2023; Shi et al., 2024), we argue that LRMs pose their own unique safety challenges that require dedicated analysis. This paper aims to bridge this gap by providing a comprehensive examination of safety considerations specific to reasoning-enhanced models.
@@
In this survey, we begin with an introduction to the background of LRMs (Section 2). We then explore the safety risks of LRMs across various scenarios and settings (Section 3). Next, we demonstrate how LRMs are vulnerable to adversarial attacks and categorize these attacks based on their objectives (Section 4). We proceed to examine various defense strategies to mitigate these risks and attacks (Section 5). Finally, we outline promising future research directions (Section 6). A timeline depicting the evolution of different approaches is shown in Figure 1. The comprehensive structure of our survey is illustrated in Figure 2.
@@ 2 Background
for tree= font=, draw=myblue, semithick, rounded corners, minimum height = 1.ex, minimum width = 3em, anchor = west, grow = east, forked edge, s sep = 2mm, fork sep = 1mm, [Safety in LRMs: A Survey,rotate=90,anchor=center, [Defenses for LRMs (Sec. 5), fit=band, text width=1.7cm, fill=defenseblue, draw=blueborder [Guard Models for LRMs (Sec. 5.3), text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Reasoning-based Guard Model, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [GuardReasoner (Liu et al., 2025b), ThinkGuard (Wen et al., 2025), X-Guard (Upadhayay et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] [Classifier-based Guard Model, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [LLaMA Guard 3 (Dubey et al., 2024), Aegis Guard 2 (Ghosh et al., 2024b), WildGuard (Han et al., 2024), ShieldGemma (Zeng et al., 2024a), LLaMA Guard 3-Vision (Chi et al., 2024a), Beaver Guard-V (Ji et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] ] [Inference-time Defenses for LRMs (Sec. 5.2), text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Safe Decoding for Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [ZeroThink/LessThink/MoreThink (Jiang et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] [Inference-time Scaling on Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Inference-time Compute (Zaremba et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] ] [Safety Alignment of LRMs (Sec. 5.1), text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [RL-based Safety Alignment on Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Deliberative Alignment (Guan et al., 2024), STAIR (Zhang et al., 2025c), SaRO (Mou et al., 2025), R2D (Zhu et al., 2025a), text width=6cm, fill=defenseblue, draw=blueborder] ] [SFT-based Safety Alignment on Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [SafeChain (Jiang et al., 2025), RealSafe-R1 (Zhang et al., 2025b), text width=6cm, fill=defenseblue, draw=blueborder] ] [Safe CoT Data Curation, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [STAR-1 (Wang et al., 2025), SafeChain (Jiang et al., 2025), RealSafe-R1 (Zhang et al., 2025b), text width=6cm, fill=defenseblue, draw=blueborder] ] ] ] [Attacks on LRMs (Sec.4), fit=band, text width=1.7cm, fill=attackred, draw=blueborder [Jailbreak Attacks (Sec. 4.4), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Reasoning-based Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Mousetrap (Yao et al., 2025), H-CoT (Kuo et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] [Multi-Turn Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [ActorAttack (Ren et al., 2024), RACE (Ying et al., 2025a), MHJ (Li et al., 2024), text width=6cm, fill=attackred, draw=blueborder] ] [Prompt-based Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Past Tense Andriushchenko and Flammarion (2024), CNSafe Ying et al. (2025b), SafeMLRM Fang et al. (2025), text width=6cm, fill=attackred, draw=blueborder] ] ] [Prompt Injection Attacks (Sec. 4.3), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Nerd Sniping Zaremba et al. (2025), R1 Assessment Zhou et al. (2025), text width=6cm, fill=attackred, draw=blueborder] ] [Answer Correctness Attacks (Sec. 4.2), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Error Injection, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [CPT (Cui et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] [Reasoning-based Backdoor Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [BadChain (Xiang et al., 2024), DarkMind (Guo and Tourani, 2025), BoT (Zhu et al., 2025b), ShadowCoT (Zhao et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] ] [Reasoning Length Attacks (Sec.4.1), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Underthinking, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Think Less (Zaremba et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] [Overthinking, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [OverThink attack (Kumar et al., 2025), Nerd Sniping (Zaremba et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] ] ] [Safety Risks of LRMs (Sec.3), fit=band, text width=1.7cm, fill=riskyellow, draw=blueborder [Multi-modal Safety Risks (Sec. 3.4), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [SafeMLRM (Fang et al., 2025), text width=6cm, fill=riskyellow, draw=blueborder] ] [Multi-lingual Safety Risks (Sec. 3.3), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [CHiSafetyBench Zhang et al. (2025a), CNSafe Ying et al. (2025b), ALIA Romero-Arjona et al. (2025), text width=6cm, fill=riskyellow, draw=blueborder] ] [Agentic Misbehavior Risks (Sec. 3.2), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [HHH Trade-offs Xu et al. (2025), Medical Cyberattack Qiu et al. (2025), Specification Gaming Bondarenko et al. (2025), Self-preservation Barkur et al. (2025), InstrumentalEval He et al. (2025), text width=6cm, fill=riskyellow, draw=blueborder] ] [Harmful Request Compliance Risks (Sec.3.1), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [DeepSeek Thoughtology Marjanović et al. (2025), ASTRAL Arrieta et al. (2025a), o3vsDeepSeek Arrieta et al. (2025b), text width=6cm, fill=riskyellow, draw=blueborder] ] ] ]
@@ 3 Safety Risks of LRMs
As LRMs continue to advance, they introduce distinct safety challenges that warrant careful examination even in standard, non-adversarial scenarios. The explicit reasoning processes that make these models powerful become potential vectors for harm during routine operation. In this section, we examine four key categories of inherent safety risks: unsafe request compliance (Section 3.1), multi-lingual safety disparities (Section 3.3), concerning agentic behaviors (Section 3.2), and multi-modal safety challenges (Section 3.4). Understanding these fundamental vulnerabilities is essential for developing effective safeguards and ensuring the responsible deployment of reasoning-enhanced AI systems, complementing the study of deliberate exploitation methods addressed later.
@@ 4.1 Reasoning Length Attacks
The success of step-by-step reasoning in LRMs has significantly enhanced their problem-solving capabilities, but this improvement comes with a critical vulnerability: overthinking. Recent work by Chen et al. (2024) has identified that these models often spend orders of magnitude more computation on simple questions with minimal benefit, creating substantial inference overhead and latency issues. Hashemi et al. (2025) systematically demonstrate this inefficiency through their DNR benchmark, revealing that reasoning models generate up to 70× more tokens than necessary and often perform worse than simpler non-reasoning models on straightforward tasks. This inefficiency creates an exploitable attack surface where adversaries can deliberately trigger excessive reasoning through carefully crafted inputs. Kumar et al. (2025) formalize this as an indirect prompt injection attack that introduces computationally demanding decoy problems, while Zaremba et al. (2025) identify Nerd Sniping attacks that trap models in unproductive thinking loops, causing them to spend abnormally large amounts of inference-time compute with decreased performance. These attacks effectively apply denial-of-service techniques (Shumailov et al., 2021; Gao et al., 2024) specifically to LRMs. The implications extend beyond computational waste—Marjanović et al. (2025) and Wu et al. (2025) demonstrate that reasoning performance actually degrades beyond certain length thresholds, while Cuadron et al. (2025) show that in agentic systems, overthinking can lead to decision paralysis and ineffective action selection.
@@ 4.3 Prompt Injection Attacks
Prompt injection attacks affect both traditional LLMs and LRMs, but LRMs present distinct challenges due to their step-by-step processing. These attacks (Kumar et al., 2024; Liu et al., 2023) inject malicious instructions disguised as normal user input, causing the AI to override or ignore its original developer-set instructions and safeguards. The explicit reasoning structures of LRMs offer attackers additional insertion points to redirect the model’s thought process, potentially making them more susceptible to certain types of injections.
@@ 5.1 Safety Alignment of LRMs
First, Wang et al. (2025) build a 1k-scale safety dataset named STAR-1 specifically designed for LRMs. Another safety training data in CoT style named SafeChain (Jiang et al., 2025) is introduced to enhance the safety of LRMs. In addition, Zhang et al. (2025b) construct a dataset consisting of 15k safety-aware reasoning trajectories, generated by DeepSeek-R1, with explicit instructions designed to promote expected refusal behavior.
@@
Based on the curated safe CoT data, researchers further conduct SFT to improve safety. For example, Jiang et al. (2025) train two LRMs with the SafeChain dataset, demonstrating that it not only enhances model safety but also preserves reasoning performance. Besides, RealSafe-R1 (Zhang et al., 2025b) is developed to make LRMs safer by training DeepSeek-R1 distilled models on the 15k safety-aware reasoning trajectories.

### v3

@@ Abstract
Large Reasoning Models (LRMs) have exhibited extraordinary prowess in tasks like mathematics and coding, leveraging their advanced reasoning capabilities. Nevertheless, as these capabilities progress, significant concerns regarding their vulnerabilities and safety have arisen, which can pose challenges to their deployment and application in real-world settings. This paper presents the first comprehensive survey of LRMs, meticulously exploring and summarizing the newly emerged safety risks, attacks, and defense strategies specific to these powerful reasoning-enhanced models. By organizing these elements into a detailed taxonomy, this work aims to offer a clear and structured understanding of the current safety landscape of LRMs, facilitating future research and development to enhance the security and reliability of these powerful models.
@@ 1 Introduction
Large Language Models (LLMs) (Meta, 2024; Qwen et al., 2025; Ke et al., 2025) have achieved remarkable proficiency across tasks ranging from open-domain conversation to program synthesis. Central to their utility is reasoning: the ability to derive logically coherent conclusions by chaining together intermediate inferences.
@@
The recent release of OpenAI’s o1 series (OpenAI, 2024) marks the emergence of Large Reasoning Models (LRMs) (Li et al., 2025b), which are explicitly trained to produce richly formatted, human-readable reasoning traces. Notable examples include DeepSeek-R1 (DeepSeek-AI et al., 2025), Kimi-1.5 (Team et al., 2025), and QwQ (Team, 2024b), all of which leverage reinforcement learning to refine their deduction processes. LRMs now set new benchmarks in mathematical problem solving (Lightman et al., 2023), closed-book question answering (Rein et al., 2024), and code generation (Jain et al., 2024).
As LRMs become increasingly integrated into high-stakes domains—from scientific research to autonomous decision support—it is vital to rigorously assess their safety, robustness, and alignment. Despite the existence of surveys on LLM safety (Huang et al., 2023; Shi et al., 2024), the enhanced capabilities of LRMs make it important to perform a dedicated analysis of their unique safety challenges. This paper aims to bridge this gap by providing a comprehensive examination of safety considerations specific to reasoning-enhanced models.
@@
In this survey, we begin with an introduction to the background of LRMs (Sec. 2). We then explore the inherent safety risks of LRMs (Sec. 3), focusing on vulnerabilities that emerge during standard, non-adversarial usage scenarios without deliberate exploitation attempts. Next, we distinguish these inherent risks from deliberate adversarial attacks (Sec. 4), where we categorize and analyze techniques specifically designed to compromise or manipulate LRMs’ reasoning capabilities. We proceed to examine various defense strategies to mitigate both inherent risks and adversarial attacks (Sec. 5). Finally, we outline promising future research directions (Sec. 6). A timeline depicting the evolution of different approaches is shown in Figure 1. The comprehensive structure of our survey is illustrated in Figure 2.
@@ 2 Background
for tree= font=, draw=myblue, semithick, rounded corners, minimum height = 1.ex, minimum width = 3em, anchor = west, grow = east, forked edge, s sep = 2mm, fork sep = 1mm, [Safety in LRMs: A Survey,rotate=90,anchor=center, [Defenses for LRMs (Sec. 5), fit=band, text width=1.7cm, fill=defenseblue, draw=blueborder [Guard Models for LRMs (Sec. 5.3), text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Reasoning-based Guard Model, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [GuardReasoner (Liu et al., 2025b), GuardReasoner-VL, (Liu et al., 2025d), ThinkGuard (Wen et al., 2025), X-Guard (Upadhayay et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] [Classifier-based Guard Model, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [LLaMA Guard 3 (Dubey et al., 2024), Aegis Guard 2 (Ghosh et al., 2024b), WildGuard (Han et al., 2024), ShieldGemma (Zeng et al., 2024a), LLaMA Guard 3-Vision (Chi et al., 2024a), Beaver Guard-V (Ji et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] ] [Inference-time Defenses for LRMs (Sec. 5.2), text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Safe Decoding for Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [ZeroThink/LessThink/MoreThink (Jiang et al., 2025), Thinking Intervention (Wu et al., 2025a), text width=6cm, fill=defenseblue, draw=blueborder] ] [Inference-time Scaling on Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Inference-time Compute (Zaremba et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] ] [Safety Alignment of LRMs (Sec. 5.1), text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [RL-based Safety Alignment on Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [Deliberative Alignment (Guan et al., 2024), STAIR (Zhang et al., 2025c), SaRO (Mou et al., 2025), R2D (Zhu et al., 2025a), text width=6cm, fill=defenseblue, draw=blueborder] ] [SFT-based Safety Alignment on Reasoning, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [SafeChain (Jiang et al., 2025), RealSafe-R1 (Zhang et al., 2025b), RT (Wang et al., 2025a), Safety Tax (Huang et al., 2025), text width=6cm, fill=defenseblue, draw=blueborder] ] [Safe CoT Data Curation, text width=2.5cm, l sep = 2mm, fill=defenseblue, draw=blueborder [STAR-1 (Wang et al., 2025b), SafeChain (Jiang et al., 2025), RealSafe-R1 (Zhang et al., 2025b), text width=6cm, fill=defenseblue, draw=blueborder] ] ] ] [Attacks on LRMs (Sec.4), fit=band, text width=1.7cm, fill=attackred, draw=blueborder [Jailbreak Attacks (Sec. 4.4), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Reasoning-based Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Mousetrap (Yao et al., 2025), H-CoT (Kuo et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] [Multi-Turn Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [ActorAttack (Ren et al., 2024), RACE (Ying et al., 2025a), MHJ (Li et al., 2024), text width=6cm, fill=attackred, draw=blueborder] ] [Prompt-based Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Past Tense Andriushchenko and Flammarion (2024), CNSafe Ying et al. (2025b), SafeMLRM Fang et al. (2025), text width=6cm, fill=attackred, draw=blueborder] ] ] [Prompt Injection Attacks (Sec. 4.3), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Nerd Sniping Zaremba et al. (2025), R1 Assessment Zhou et al. (2025), text width=6cm, fill=attackred, draw=blueborder] ] [Answer Correctness Attacks (Sec. 4.2), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Error Injection, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [CPT (Cui et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] [Reasoning-based Backdoor Attacks, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [BadChain (Xiang et al., 2024), DarkMind (Guo and Tourani, 2025), BoT (Zhu et al., 2025b), ShadowCoT (Zhao et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] ] [Reasoning Length Attacks (Sec.4.1), text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Underthinking, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [Think Less (Zaremba et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] [Overthinking, text width=2.5cm, l sep = 2mm, fill=attackred, draw=blueborder [OverThink attack (Kumar et al., 2025), Nerd Sniping (Zaremba et al., 2025), text width=6cm, fill=attackred, draw=blueborder] ] ] ] [Safety Risks of LRMs (Sec.3), fit=band, text width=1.7cm, fill=riskyellow, draw=blueborder [Multi-modal Safety Risks (Sec. 3.4), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [SafeMLRM (Fang et al., 2025), text width=6cm, fill=riskyellow, draw=blueborder] ] [Multi-lingual Safety Risks (Sec. 3.3), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [CHiSafetyBench Zhang et al. (2025a), CNSafe Ying et al. (2025b), ALIA Romero-Arjona et al. (2025), text width=6cm, fill=riskyellow, draw=blueborder] ] [Agentic Misbehavior Risks (Sec. 3.2), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [HHH Trade-offs Xu et al. (2025), Medical Cyberattack Qiu et al. (2025), Specification Gaming Bondarenko et al. (2025), Self-preservation Barkur et al. (2025), InstrumentalEval He et al. (2025), text width=6cm, fill=riskyellow, draw=blueborder] ] [Harmful Request Compliance Risks (Sec.3.1), text width=2.5cm, l sep = 2mm, fill=riskyellow, draw=blueborder [DeepSeek Thoughtology Marjanović et al. (2025), ASTRAL Arrieta et al. (2025a), o3vsDeepSeek Arrieta et al. (2025b), text width=6cm, fill=riskyellow, draw=blueborder] ] ] ]
@@ 3 Safety Risks of LRMs
As LRMs continue to advance, they introduce distinct safety challenges that warrant careful examination even in standard, non-adversarial scenarios. The explicit reasoning processes that make these models powerful become potential vectors for harm during routine operation. In this section, we examine four key categories of inherent safety risks: unsafe request compliance (Sec. 3.1), multi-lingual safety disparities (Sec. 3.3), concerning agentic behaviors (Sec. 3.2), and multi-modal safety challenges (Sec. 3.4). Understanding these fundamental vulnerabilities is essential for developing effective safeguards and ensuring the responsible deployment of reasoning-enhanced AI systems, complementing the study of deliberate exploitation methods addressed later.
@@ 4.1 Reasoning Length Attacks
The success of step-by-step reasoning in LRMs has significantly enhanced their problem-solving capabilities, but this improvement comes with a critical vulnerability: overthinking. Recent work by Chen et al. (2024a) has identified that these models often spend orders of magnitude more computation on simple questions with minimal benefit, creating substantial inference overhead and latency issues. Hashemi et al. (2025) systematically demonstrate this inefficiency through their DNR benchmark, revealing that reasoning models generate up to 70× more tokens than necessary and often perform worse than simpler non-reasoning models on straightforward tasks. This inefficiency creates an exploitable attack surface where adversaries can deliberately trigger excessive reasoning through carefully crafted inputs. Kumar et al. (2025) formalize this as an indirect prompt injection attack that introduces computationally demanding decoy problems, while Zaremba et al. (2025) identify Nerd Sniping attacks that trap models in unproductive thinking loops, causing them to spend abnormally large amounts of inference-time compute with decreased performance. These attacks effectively apply denial-of-service techniques (Shumailov et al., 2021; Gao et al., 2024) specifically to LRMs. The implications extend beyond computational waste—Marjanović et al. (2025) and Wu et al. (2025b) demonstrate that reasoning performance actually degrades beyond certain length thresholds, while Cuadron et al. (2025) show that in agentic systems, overthinking can lead to decision paralysis and ineffective action selection.
@@ 4.3 Prompt Injection Attacks
Prompt injection attacks affect both traditional LLMs and LRMs, but LRMs present distinct challenges due to their step-by-step processing. These attacks (Kumar et al., 2024; Liu et al., 2023; Chen et al., 2024b, 2025) inject malicious instructions disguised as normal user input, causing the AI to override or ignore its original developer-set instructions and safeguards. The explicit reasoning structures of LRMs offer attackers additional insertion points to redirect the model’s thought process, potentially making them more susceptible to certain types of injections.
@@ 5.1 Safety Alignment of LRMs
First, Wang et al. (2025b) build a 1k-scale safety dataset named STAR-1 specifically designed for LRMs. Another safety training data in CoT style named SafeChain (Jiang et al., 2025) is introduced to enhance the safety of LRMs. In addition, Zhang et al. (2025b) construct a dataset consisting of 15k safety-aware reasoning trajectories, generated by DeepSeek-R1, with explicit instructions designed to promote expected refusal behavior.
@@
Based on the curated safe CoT data, researchers further conduct SFT to improve safety. For example, Jiang et al. (2025) train two LRMs with the SafeChain dataset, demonstrating that it not only enhances model safety but also preserves reasoning performance. Besides, RealSafe-R1 (Zhang et al., 2025b) is developed to make LRMs safer by training DeepSeek-R1 distilled models on the 15k safety-aware reasoning trajectories. Wang et al. (2025a) proposes training the model to reason with the guidelines, thereby enhancing survey alignment.